|Current Location||Sensorica Main Lab|
Power supply GQ
Any SENSORICA affiliate, Fablab member of Techshop member.
Custodian : SENSORICA
Manager : Tibi (Contact)
|Record created date||Feb. 3, 2013|
|Record created by||None|
We don't know.
Design and prototype the Water producing system.
Design and prototype the light supplement system for the iPot
Preliminary experiments with the Micro 3D printer.
Preliminary experiments with the Micro 3D printer.
creating prototype for PMMA joint-type transducer.
prototype the low cost tape sensor, will be taken by Matheus
Worked with Jonathan, figuring out how to use Frederic's electronics for the Mosquito SS4
NOTE: I need to log Frederic's electronics in the system!!
3-Axis piezo driver assembled and working.
The driver is on operation with the piezo stack and the piezo tube for 1h for test. The metallic enclosure is warm, but just a little, nothing to be aware.
New components added to the part list.
Re-assembling a 3-Axis driver using the new PCB and updating the part list with the missing components
Connectorizing the 3-axis piezo finger. Instructions here: https://docs.google.com/document/d/1qGbzFsd7eTMJ2wGnYq7mnx2ik7y9CZoJWg2onIih7BY
Proceeding tests, The electronic system is working well, but the frequency limit is lower than we've supposed (about 25Hz using the filtered output, accordint to the datasheet we expect 3x this frequency using the non-filtered output).
The fiber is moving on the Z-axis, but we have a XY component working at the same time. I think this is due the bad quality of the actual mechanics and the non uniformity of the springs. I've also found a mistake at the top_base, it should the rotated by 45 degrees to be aligned with the XY-axis!
Proceeding a heating test on the 3-axis PDm200: (https://docs.google.com/spreadsheet/ccc?key=0AvKN6G919YgndEEyamtKX1ZFU0Zkb1pIUTdFS2hybWc&usp=drive_web#gid=4)
1) no external dissipation, box closed, sensor on the black heat sink, measuring each 5min
2) Using the thermal pad with adhesive (H48-6G-20-20-2.6-1A) and NO thermal paste connecting the PDm200 to the aluminum box
No load or input
3) Using the thermal pad with adhesive (H48-6G-20-20-2.6-1A) and DO using thermal paste connecting the PDm200 to the aluminum box
No load or input
The tests showed the need of using the Thermal Pad but the use of Thermal Paste have not increase the dissipation capacity. A second test without thermal paste didn't have the same result, probable because the coupling between the 2 surfaces, a thicker thermal pad should solve this problem. The thermal pad data sheet shows that some pressure is necessary to obtain an optimal response.
Redesign base PCB to fit 10cm long specification and reallocate components:
Design the Front and End plates:
K End Plate.SLDDRW
K Front Plate.SLDDRW
PDF versions saved here: https://drive.google.com/?tab=mo&authuser=0#folders/0B_KN6G919YgnTmRpNVQyYU1XR1U
search of components: BNC cable, piezo buzzer, thermal adhesive pad
R&D work on the Piezo Driver
panelizing pcbs and making order
regenerating gerbers with separate slot drill file
Finalized pcb and generated gerber files and drill files.
base board rev2 is now finished, silk overlay is finished
1/7/2014 modifying Piezo Drive Module Integration base board - rev 2 which can be found here:
\Dropbox\tactus\piezo system\PCB - rev 4 - Piezo Drive Module integration base board - rev2\Piezo Drive Module Integration base board - rev 2.PrjPcb
I am adjusting the pcb to match the dimensions given by antonio in this document:
Piezo Driver Connector search, quoted PDu100 modules from american supplier 1.5hrs
Piezo PCB and component search 10h55mins ~11h
Piezo PCB still on being worked on. Today spent 3hrs and will press start button now.
See pcb here:
\Dropbox\tactus\piezo system\PCB - rev 4 - Piezo Drive Module integration
Piezo Module PCB Layout 40m
Piezo Module PCB Layout 6h05
Piezo Module PCB Layout 4h30
Piezo Module PCB Layout 2h52
Piezo Module PCB Layout 4h50
Piezo Module PCB Layout 4h30
piezo pdm200 testing, sent po for linear stage 7h50
Piezo Module PCB layout + meeting 3h25
Piezo Clean up (sorted all the accumulated parts, sorted work spaces etc) 4h30
Piezo Module PCB layout 3h45
Baseboard PCB layout 6h
Baseboard PCB layout 2h
Piezo 2ch Assembly 2h50
3Ch piezo final design review, start soldering 2ch piezo 3h30
Designing OUT board pcb 3h30
See it here \Dropbox\tactus\piezo system\PCB - rev 4 - Piezo Drive Module integration OUT board
Designing PCB 8h25
Testing 3 Ch Piezo and designing PCB 5h18
Designing 3ch PCB 3h
see it here: \Dropbox\tactus\piezo system\PCB - rev 4 - Piezo Drive Module integration base board
Patterning and Etching PCBs for 3Ch board and 2ch out board 8h
Cutting PCBs 45m
PCB Assembly 26m
Assembling Piezo PCBs 26m
Above three items are for the home made pcbs, the out board and the base board, for the 3 ch device
Assembling Piezo PCBs 3h35
Recording the time it takes to assemble piezo pcbs 4h45
see document here: https://docs.google.com/a/olesik.com/document/d/1qGbzFsd7eTMJ2wGnYq7mnx2ik7y9CZoJWg2onIih7BY/edit
Ordering Piezo Parts (for extra units) 4h10
This is for the Piezo Prototypes created by F.Cantin
piezo bom and layout 5h
piezo layout 7h45
piezo box shopping, final piezo layout 8h20
piezo bom 3h20
pcb search and order 1h40
Working on the interface for the piezo 1h
pcb payout for adapter board 20m
piezo assembly 4h40
testing piezo 6h
piezo debugging 5h15
pcb piezo prototype 3h20
pcb piezo prototype 1h
etching piezo prototype 5h
soldering piezo 6h30
June 17, 2013
See documentation in this doc
The setup was made before, there is another labnote for it.
June 17, 2013
Continued the work. See the Google doc for more details. Jonathan and Antonio were also involved.
We discovered that the analog out of the Labjack is limited to steps of 0.02Volts. We need 0.002Volts resolution, in order to test below 0.5um piezo steps. Jonathan will make a circuit for this.
July 18, 2013
Continued work on characterization. I am doing long acquisitions for precision tests. These results will be entered in the document in the Precision section.
August 01, 2013
Worked with Bing and Antonio on the Piezo. We have a problem with the assembly of the piezo controller. Jonathan made a new prototype and we tested it and it did not work. Frederic's prototype still works fine. We could not clearly understand why the second prototype did not work. The new boards for the product seam to be fine. So the problem is still a mystery.
August 05, 2013
Recreated the characterization setup, because Jonathan had taken the piezo driver prototype away for the fabrication. I modified the way the fiber is attached to the piezo stack.
August 06, 2913
Resumed characterization experiments. I worked on Precision. See more
The long-term stability problem prevents us from directly measure precision. I put a note in the doc and sent message to the team about it.
"This [long-term stability of the Mosquito] is a problem with the acquisition system and with the Mosquito in general that we need to address! We need to improve the architecture of the Mosquito by integrating a reference, which is divided from the signal to account for intensity fluctuations."
I entered the best data here.
One way around the stability is to measure the difference between 0 Volt and the x Volt, "x" representing an input voltage to the piezo controller, which will result in a motion step.
Data still needs to be processed.
Oct 23, 2013,
tested the 2 axis piezo system, driver and the actuator. Worked with Antonio and Jonathan. Produced the piezo manual
Other work was logged previously as non-production work.
The goal is to build a setup and a protocol to test and characterize the xyz piezo.
The deliverable is a report and piezo specs sheet.
Piezo driver tested
CREATE DESIGN AND PROTOTYPE AND ENTER AS INPUT
Piezo tested is a piezo stack - PROVENANCE
The previous days I mounted an optical fiber stretch sensor on the Piezo.
I also prepared the LabView program for the tests.
JULY 03, 2013
Today I worked with Jonathan on mounting the new LED 850nm Mosquito to monitor the piezo motion.
We tried to use the Chinese MM circulator in the Mosquito, instead of the Y coupler/splitter, but no light seems to pass through it. the LED and the PD/TIA were coupled directly to make sure that they functioned properly.
I tried to make a Y coupler/splitter using our microsplicer - fusing 3 fibers together, but I was unsuccessful in the first trial. The second one was not completed today, testing will be done tomorrow.
Jul 10 2013
I had to redo the optical fiber attachment to the piezo. The problem was that the piezo was not connected to the fiber holders and was sliding under protective tape. I had to directly connect the piezo the both ends of the piezo ceramic. This her arrangement was tested succesfully. I also adjusted the gain of the mosquito to increase the swing of the signal. We are using an LED mosquito at 850 nm, with the circuit of a mosquito prototype made by Jonathan for the 1550 nm mosquito. I am also using Frederic D laser driver which I connected to a power supply.
This is for Ronan
These activies are repeated in order to clarify the procedures of the Low-cost Tape Sensor design. The citations include these steps as well.
Tape sensor - introduction (done)
Tape sensor - gap optimization (IN PROGRESS)
Tape and flex sensor - comparison (IN PROGRESS)
Tape sensor - 3D printing (IN PROGRESS)
Tape sensor - manufacturing (STANDBY)
Created a low cost optical tape sensor that can be put on different structures to detect bending.
We'll apply it first on a hockey stick.
See project page
Started on Friday 27, 2012 After we created the prototype and made the demo for Zhu (Canadian Space Agency, beam deflection sensor) we realized that this device works in stretching and compression. I had the idea to create a very simple and low cost tape sensor using 250 um diameter PMMA fiber and kapton tape. The goal of this experiment was to build the first such prototype and test it. See design https://docs.google.com/drawings/d/17Wb68vkZPeV5jz1GJDJG-TXER2Ofy-1wyndfQbkgp10/edit The device was built and successively tested. See the video made with Daniel http://youtu.be/yKvdryt1iKk
04 JUNE 2013
Worked on the setup to fabricate the low cost Tape transducer. I continued in the same direction from where I left it. The device is now almost done, need to take some pictures and document how it works. This is still a prototype device, to produce a few of them in a systematic matter and test them.
I will use the enhanced LED 850nm Mosquito to test it. I need to make a project presentation to SENSORICA about this, in a week.
05 JUNE 2013
Did some mechanical work on the fabrication device. I finally put in place the micrometer screw and 2 springs to move the working plate. It seems to be solid. After this, I shifted my attention to the Y connector for the 250um diameter PMMA fiber. I created a separate labnotes where I marked 7 hours of work on this. The main problem was to create a mold/stamp for PCL, from which I make the connector. The idea is to stamp the Y grove into PCL, place the fibers into the grooves and apply optical clear epoxy to hold the fibers aligned in place. I started with clay, and moved to etching aluminum.
10 JUNE 2013
Produced a new low-cost tape sensor using Francois' design.
11 JUNE 2013
work on the tape sensor presentation
14 JUNE 2013
Created documentation for the low cost tape sensor, adding stuff on the website, and updating the document.
Gave training to Ronan on the Tape Sensor. I also created a labnote for him to collaborate on this project.
16 AUGUST 16, 2013
Worked with Rodrigo on a new tape sensor design. We are trying to implement the idea of optical gearing. See concepts here
One design use a concave lens (perhaps made with PDMS, using rounded lass rod/fiber as mold). See idea here
The other one uses a scattering (+ absorbent) medium within the gap between two fibers, using index matching gel with micron-size particles (silica powder or carbon powder).
See the design for the second choice here.
Rodrigo and I designed an experiment to compere intensity variation with gap between an air gap and a scattering (+absorbent) medium. We are going to create the setup and perform the comparison next week. The deliverable will be a white paper. See idea here.
23 AUGUST 2013
Worked on the low cost tape sensor with Rodrigo, helped him with the optical setup for some exploratory testing of optical gears. We are using 1mm PMMA fiber for these tests. This is in continuation of the previous work.
2013 May 21 (5.5 h + 1.5 h):
prepare 5 joint-type transducers in angle using Ivan's pre-bent glass capillaries. Result: broke one, fried one, one has a damaged mirror. 2 left to connectorize. Connectorization failed. 0 transducer made....
2013 May 22 (4h):
goal: mirror 4 levers, connectorize one transducer.
silvering quality tests: played with number of silver layers and reaction time:
red lever: 3 times one minutes shaking. Intensity = ?
one black strip lever: one time one minute shaking. Intensity = ?
two black strips lever: one time 2 minutes. Intensity = ?
Arranged power supply memories for shrinking:
M1 = 0
M2 = 0.32 A x 0.67 V = 0.2 W
M3 = 0.4 x 1 = 0.4 W
M4 = 0.6 x 0.1 = 0.6 W
Determined ideal length for delivery fiber at the very beginning (before cleaving 1 cm both side) = 16.7 cm
Result: made one pre-bent joint-type transducer.
2013 May 24 (0.5 h):
tested transducer with LED Mosquito: it works
2013 June 5 (7.5 h)
goals: test new LED-Mosquito with amplification + filter
1) signal tests on LED-Mosquito: signal without anything plugged (no delivery fiber): 0.995-1.065 V (average 1030 mV), noise = 70m V.
+ delivery fiber: 1.205-1.125 V = 80mV noise (average baseline 1165 mV), so the delivery fiber increases reflexion of 135 mV.
Reference: with industrial mirror max signal goes to + 10.3 V (note that signal goes up now when light goes back to the PD, it was the opposite before: signal was going down when more light).
Note 2: signal saturates when lever with mirror in front of delivery fiber, so gain is too high.
2) Shrinking trial 1: OK, Shrinking trial 2: failed: shrinking tube didn't stick to the fibers, dust or grease ? Didn't stick even with washing with alcohol.
3) connectorization procedure update:
step 1) verify if chinese ST connector is not clogged. Place the conenctor horizontally on the table and push a stripped fiber in. Don't maintain the connector so if the fiber can't go inside, the fiber won't break, it will just oush the connector.
step 2) assemble the glass capillary and the black plastic cover
step 3) insert the shrunk fiber into the glass capillary
step 4) use the same technique as in step 1) to insert the shrunk fiber free end into the connector
step 5) adjust the gap position relative to the capillary end by gently sliding the capillary. Apply glue on the connector to fix the black cover. Add glue just below the gap between gap and capillary end.
2013 June 6 (1.5 h)
Worked with Francois on exploring the constriction transducer with the LED 850nm Mosquito.
Worked on the constriction again. I made some experiments to distinguish between my model and Frederic's model of the constriction transducer.
My model: some light get's transferred into the cladding at the constriction site and after some travel comes back in the core. As it does that, it interferes with the light that continued into the core and the detector sees this interference pattern. I thought that if I coat the entire constricted area and the lever with silver I would increase the sensitivity of the transducer, because more light would come back into the core.
I made 3 transducers, one with 4 constrictions, one with 2 constrictions and the one with a single constriction. The first two I constricted approx 30%, the last one approx 50%. I connected them to the Chinese LD Mosquito that came back from Phil.
Through the fluctuations of this Mosquito, because we understand its defaults, I could measure some sensitivity for the 4 constrictions transducer, but not good enough for the other ones. The Mosquito behaves in a strange/unpredictable way.
I used the reusable optical fiber connectors to connect the fiber.
In conclusion, it seems that my theory about the mechanism behind the constriction transducer is not the one I thought.
Pictures and videos were made.